OpenStudy (anonymous):

Help !! is -7+9=-9+7 true, false, or open ??

OpenStudy (solomonzelman):

false

OpenStudy (solomonzelman):

2 doesn't equal -2

OpenStudy (anonymous):

this is true -(-7 +9) = -9+7

OpenStudy (solomonzelman):

Yes, but that is a different problem.

OpenStudy (solomonzelman):

@ws18, did this help you?

OpenStudy (anonymous):

so ... its true ?

OpenStudy (solomonzelman):

NO

OpenStudy (solomonzelman):

False, -7+9=2 and -9+7=-2 See?

OpenStudy (anonymous):

ok thank you

OpenStudy (solomonzelman):

TNX